# Online Conformal Model Selection for Nonstationary Time Series

### v1

@@ 1 Introduction
The MPS is highly generic. It does not rely on the identification of a true model, nor does it require the time series to be stationary. The definition of the best model is also flexible: the process of the best models $\{\mathcal{M}_{t}\}$ can be defined based on any user-chosen empirical model evaluation metric, such as in-sample measures (e.g., information criteria, and likelihood- or residual-based diagnostic measures) or out-of-sample measures like forecast accuracy at any forecast horizon. Moreover, the model class $\mathscr{M}$ can be any statistical models or black-box machine learning algorithms. In fact, as discussed in Section 2, the term “models” can broadly refer to any models, learning or forecasting algorithms, and even alternatives such as policies that do not necessarily involve modeling the data. Additionally, $t$ may represent an individual time point at which data are collected, but more generally, it may refer to the end of the $t$th time period, with each time period encompassing multiple time points. Hence, we use the terms time period and time point interchangeably throughout this paper.
@@
Arnold et al., (2024) proposes sequential MCS by incorporating sequential testing methods, offering time-uniform, nonasymptotic coverage guarantees. However, as it is developed within the MCS framework, it is still inherently an offline method; that is, the coverage is defined for the current rather than the next time period. MCS has also been extended to various contexts, including high-dimensional settings (Dufour and Luger,, 2021) and environments with structural breaks (Rossi and Sekhposyan,, 2022). Recent works by Chernozhukov et al., (2021) and Ramdas et al., (2023) develop sequential testing frameworks for dynamic data, but these methods focus on parameter inference rather than model selection. At a high level, MPS is also related to the extensive literature on conformal prediction, which was first introduced by Vovk et al., (2005) as a model-agnostic, distribution-free framework for finite-sample uncertainty quantification, with marginal coverage guarantees under exchangeability; see also Lei et al., (2018). Recent extensions have expanded its scope to handle dependent and complex data structures, such as the work of Tibshirani et al., (2019) on covariate shifts and Zhou and Müller, (2024) on random objects. On the other hand, adaptive conformal inference introduced by Gibbs and Candès, (2021), as well as its variants such as Zaffran et al., (2022); Gibbs and Candès, (2024) and Yang et al., (2024), aims to calibrate time-varying nominal miscoverage rate to address distribution shifts in online settings. While these works focus exclusively on predicting outcome variables, the expansion from static to dynamic settings reflects the growing need for robust uncertainty quantification in real-time systems.
@@ 2 Preliminaries
We employ the following notation throughout the paper. For any positive integer $m$, let $[m]:=\{1,\dots,m\}$. For any set $\mathcal{S}$, we denote its cardinality by $|\mathcal{S}|$. The indicator function $\mathbf{1}(\cdot)$ takes the value one if the condition is true, and zero otherwise. The Euclidean norm of a vector is denoted by $\|\cdot\|_{2}$.
@@ 3.1 MPS Procedure
[…]
@@
where $\lambda_{\max}>0$ is a pre-specified threshold. Since $\lvert\mathcal{C}_{t}(1-\alpha)\rvert$ is non-increasing in $\alpha$ while $\mathbf{1}(\alpha>\beta_{t})$ is non-decreasing in $\alpha$, the trade-off between efficiency (i.e., the cardinality of the model set) and validity (i.e., control of the miscoverage rate) is balanced by $\lambda_{t}$, which is a relative weight to penalize miscoverage at time $t$. In addition, $\lambda_{t}$ is adaptively updated by $\lambda_{t+1}=\lambda_{t}+\gamma[\mathbf{1}(\alpha>\beta_{t})-\bar{\alpha}]$, where $\gamma>0$ is the step size. This update rule ensures that a miscoverage at time $t-1$ leads to a larger $\lambda_{t}$, up to a maximum threshold $\lambda_{\max}$, and is the key lever to achieve (1). Since $\alpha\in[0,1]$, the optimization in (2) can be easily implemented via one-dimensional grid search.
@@
1 Input: A collection of candidate models $\mathscr{M}$ indexed by $\{1,\dots,m\}$, target miscoverage level $\bar{\alpha}$, threshold $\lambda_{\max}$, relative step size $c$, step size $\gamma=c\lambda_{\max}$, block size $\tau$, initial training data $\{Y_{1},\dots,Y_{n}\}$ with $n\geq\tau$, model evaluation metric $\mathcal{L}$, bootstrap sample size $B$, grid $\mathcal{G}$
@@ 4.1 Simulation Experiments
We run Algorithm 1 with $\tau=100$, $\bar{\alpha}=0.2$, and $n=500$, so the remaining $T-n=1500$ time points are used to demonstrate the MPS updates. We benchmark MPS against the offline MCS procedure and display the results under the three designs in panels (a)–(c) of Figure 3.
@@ Results
An experiment involving actual model fitting is further conducted: We generate a time series $\{Y_{t}\}_{t=1}^{T}$ with $T=2000$ from $Y_{t}=0.3Y_{t-1}+\varepsilon_{t}+0.3\mathbf{1}(1\leq t\leq 1000)\varepsilon_{t-1}$, where $\varepsilon_{t}\overset{i.i.d}{\sim}N(0,1)$, and run Algorithm 1 with $n=1000$ and $\tau=500$. Here $\mathscr{M}$ consists of AR$(p)$ and MA$(q)$ models with $1\leq p,q\leq 5$, and $\mathcal{L}$ is the squared one-step-ahead forecast error (FE).

### v2

@@ 1 Introduction
The MPS is highly generic. It does not rely on the identification of a true model, nor does it require the time series to be stationary. The definition of the best model is also flexible: the process of the best models $\{\mathcal{M}_{t}\}$ can be defined based on any user-chosen empirical model evaluation metric, such as information criteria and likelihood- or residual-based diagnostic measures, or out-of-sample measures like forecast accuracy at any forecast horizon. Moreover, the model class $\mathscr{M}$ can be any statistical models or black-box machine learning algorithms. In fact, as discussed in Section 2, the term “models” can broadly refer to any models, learning or forecasting algorithms, and even alternatives such as policies that do not necessarily involve modeling the data. Additionally, $t$ may represent an individual time point at which data are collected, but more generally, it may refer to the end of the $t$th time period, with each time period encompassing multiple time points. Hence, we use the terms time period and time point interchangeably throughout this paper.
@@
MCS have been used in traditional econometric forecasting to compare out-of-sample predictive performance, including applications to electricity price and economic growth forecasting (Weron,, 2014; Fritz et al.,, 2025). More recently, MCS has been applied to the evaluation of machine learning–based forecasting models in energy, macroeconomic, and credit risk applications (Huang et al.,, 2021; Medeiros et al.,, 2021; Dumitrescu et al.,, 2022). Moreover, there has been a rapidly growing interest in uncertainty quantification of model selection which shares a similar objective with MCS. For example, Ferrari and Yang, (2015) and Zheng et al., (2019) introduced confidence sets for model selection via F-tests and likelihood-ratio tests, respectively, and Li and Jiang, (2024) focused on high-dimensional shrinkage model selection; see also the references therein as well as recent work such as Wang et al., (2025); Lewis and Battey, (2025); Kim and Ramdas, (2025); Cecil and Mentch, (2025). Recent extensions directly built upon Hansen et al., (2011) include Arnold et al., (2024), who proposed sequential MCS by incorporating sequential testing methods, and Bauer and Kazak, (2025), who extended the MCS framework by allowing model comparison to be conducted conditional on observable states or regimes. However, all aforementioned developments are inherently offline; that is, the guarantees are defined for the current rather than the next time period.
At a high level, MPS is also related to the extensive literature on conformal prediction, which was first introduced by Vovk et al., (2005) as a model-agnostic, distribution-free framework for finite-sample uncertainty quantification, with marginal coverage guarantees under exchangeability; see also Lei et al., (2018). Recent extensions have expanded its scope to handle dependent and complex data structures, such as the work of Tibshirani et al., (2019) on covariate shifts and Zhou and Müller, (2024) on random objects. On the other hand, adaptive conformal inference introduced by Gibbs and Candès, (2021), as well as its variants such as Zaffran et al., (2022); Gibbs and Candès, (2024) and Yang et al., (2024), aims to calibrate time-varying nominal miscoverage rate to address distribution shifts in online settings. While these works focus exclusively on predicting outcome variables, the expansion from static to dynamic settings reflects the growing need for robust uncertainty quantification in real-time systems.
@@ 2 Preliminaries
We employ the following notation throughout the paper. For any positive integer $m$, let $m]:=\{1,\dots,m\}$. For any set $\mathcal{S}$, we denote its cardinality by $|\mathcal{S}|$. The indicator function $\mathbf{1}(\cdot)$ takes the value one if the condition is true, and zero otherwise. The Euclidean norm of a vector is denoted by $\|\cdot\|_{2}$.
@@ 3.1 MPS Procedure
$\alpha_{t}^{*}=\underset{\alpha}{\min}\,\mathbb{E}_{\beta_{t}\sim F_{t}}\left\{\lvert\mathcal{C}_{t}(1-\alpha)\rvert+\lambda_{t}\max[\mathbf{1}(\alpha>\beta_{t})-\bar{\alpha},0]\right\}$ (2)
@@
where $\lambda_{\max}>0$ is a pre-specified threshold. Since $\lvert\mathcal{C}_{t}(1-\alpha)\rvert$ is non-increasing in $\alpha$ while $\mathbf{1}(\alpha>\beta_{t})$ is non-decreasing in $\alpha$, the trade-off between efficiency (i.e., the cardinality of the model set) and validity (i.e., control of the miscoverage rate) is balanced by $\lambda_{t}$, which is a relative weight to penalize miscoverage at time $t$. In addition, $\lambda_{t}$ is adaptively updated by $\lambda_{t+1}=\lambda_{t}+\gamma[\mathbf{1}(\alpha>\beta_{t})-\bar{\alpha}]$, where $\gamma>0$ is the step size. This update rule ensures that a miscoverage at time t1t-1 leads to a larger $\lambda_{t}$, up to a maximum threshold $\lambda_{\max}$, and is the key lever to achieve (1). Since $\alpha\in[0,1]$, the optimization in (2) can be easily implemented via one-dimensional grid search.
@@
1Input: A collection of candidate models $\mathscr{M}$ indexed by $\{1,\dots,m\}$, target miscoverage level $\bar{\alpha}$, threshold $\lambda_{\max}$, relative step size $c$, step size $\gamma=c\lambda_{\max}$, block size $\tau$, initial training data $\{Y_{1},\dots,Y_{n}\}$ with $n\geq\tau$, model evaluation metric $\mathcal{L}$, bootstrap sample size $B$, grid $\mathcal{G}$
@@ 4.1 Simulation Experiments
We run Algorithm 1 with $\tau=100$, $\bar{\alpha}=0.2$, and $n=500$, so the remaining Tn=1500T-n=1500 time points are used to demonstrate the MPS updates. We benchmark MPS against the offline MCS procedure and display the results under the three designs in panels (a)–(c) of Figure 3.
@@ Results
An experiment involving actual model fitting is further conducted: We generate a time series $\{Y_{t}\}_{t=1}^{T}$ with $T=2000$ from $Y_{t}=0.3Y_{t-1}+\varepsilon_{t}+0.3\mathbf{1}(1\leq t\leq 1000)\varepsilon_{t-1}$, where $\varepsilon_{t}\overset{i.i.d}{\sim}N(0,1)$, and run Algorithm 1 with $n=1000$ and $\tau=500$. Here $\mathscr{M}$ consists of AR(p)(p) and MA(q)(q) models with $1\leq p,q\leq 5$, and $\mathcal{L}$ is the squared one-step-ahead forecast error (FE).